# Flow Matching at Scale: A Machine Learning Framework for Efficient Large-Size Sampling of Many-Body Systems

### v3

@@ Abstract
We propose a machine learning framework based on Flow Matching to overcome the scaling limitations of Markov Chain Monte Carlo (MCMC) methods. We demonstrate its capability in the 2D XY model, where a single network, trained only on configurations from a small ($32\times 32$) lattice at sparse temperature points, generates reliable samples for a significantly larger system ($128\times 128$) across a continuous temperature range without retraining. The generated configurations show strong agreement with key thermodynamic observables and correctly capture the signatures of the Berezinskii-Kosterlitz-Thouless (BKT) transition. This dual generalization is enabled by the Flow Matching framework, which allows us to learn a continuous, temperature-conditioned mapping. At the same time, the inductive biases of the underlying CNN architecture ensure that the learned local physical rules are scale-invariant. This ”train-small, generate-large” capability offers a powerful and efficient alternative for studying critical phenomena. The method can be directly applied to other classical or quantum many-body systems described by continuous fields on a lattice. Furthermore, this framework can serve as a powerful proposal generator in a hybrid scheme with MCMC, dramatically accelerating high-precision studies of the thermodynamic limit.
@@ I Introduction
Monte Carlo—especially Markov chain Monte Carlo (MCMC)—is one of the most powerful computational tools across three fronts: in condensed matter, it is the workhorse for probing phase transitions and complex systems [1]; in quantum many-body physics, quantum Monte Carlo is among the few reliable non-perturbative methods [2, 3]; and in lattice QCD, Euclidean-lattice Monte Carlo provides the only systematically improvable first-principles non-perturbative approach at low energies [4, 5]. Algorithmically, MCMC targets the Boltzmann measure via detailed balance and ergodicity (e.g. Metropolis, HMC [6], cluster [7]), with uncertainties governed by the integrated autocorrelation time [1].
@@
However, MCMC encounters critical slowing down as correlation lengths grow near continuous phase transitions and toward the continuum limit; in lattice QCD, topological modes can freeze, inflating autocorrelations and complicating error estimates [8]. Fermionic models further face the sign problem, whose generic resolution is NP-hard [9]. Progress relies on advanced updates [7], tempering/multilevel schemes [10], and ML-assisted samplers; notably, flow-based generative sampling have shown decorrelated proposals and efficiency gains in lattice field theories [11, 12, 13, 14, 15].
@@
Figure 1: Schematic diagram of our work. Using Flow Matching with U-Net [16] inductive biases, we learn a deterministic probability-flow sampler that—trained only on $L_{0}\times L_{0}$ lattices—simultaneously interpolates across temperatures and extrapolates (dashed lines) to physically consistent configurations of $L_{1}\times L_{1}$ lattices with $L_{1}\gg L_{0}$, revealing scale-invariant rules and enabling a fast and stable exploration of critical phenomena near the critical temperature, $T_{c}$. Here $\chi$ denotes the susceptibility, and the green shadow represents the critical regime.
@@
Early successes with models like Normalizing Flows (NFs) demonstrated the potential to eliminate critical slowing down for a single set of system parameters [11, 17]. However, the inefficiency of these initial approaches in exploring parameter space is a significant limitation. Studying critical phenomena requires simulations across various parameters (e.g., temperatures), but training a separate, computationally expensive model for each point is impractical [14]. A key advancement to address this has been the development of conditional models, which learn the system’s behavior across a continuous parameter range from sparse training data, enabling efficient interpolation and extrapolation [14, 18]. These models are often implemented with either architecturally rigid NFs, which can suffer from training instabilities like mode collapse [17, 19], or with Diffusion Models (DM) [20, 21, 22, 23, 24], which, while robust, typically rely on a computationally intensive iterative sampling process.
@@
Even as conditional models begin to address the challenge of exploring parameter space, a more fundamental and computationally demanding limitation persists in both traditional MCMC and these new generative approaches: the lack of generalizability across system sizes. Systematic studies of critical phenomena and the thermodynamic limit hinge on a sequence of simulations on progressively larger lattices. Under the current paradigm, each lattice size requires an entirely new, independent simulation or model training, a process whose computational cost scales prohibitively. Therefore, a truly efficient framework would not only interpolate across parameters like temperature but also learn the underlying, scale-invariant physical rules, enabling extrapolation from a single, small-system training process to much larger lattices. Our work is designed to address this critical gap directly.
@@
In this work, we employ Flow Matching (FM)  [25], a next-generation framework that combines the advantages of these preceding approaches while mitigating their drawbacks. FM directly learns the deterministic vector field of a continuous transformation, known as the probability flow ODE [26]. This provides a more stable regression-based training objective than traditional NFs and enables significantly faster sampling than the stochastic processes of DMs [27, 26]. This training stability, architectural flexibility, and sampling efficiency combination makes FM an ideal candidate for building a robust and generalizable conditional sampler. When integrated with the strong inductive biases of a U-Net architecture [16], we demonstrate that this superior framework solves a crucial remaining challenge: achieving a robust dual generalization. We show that a single model, trained only on data from a small lattice ($32\times 32$), can simultaneously interpolate across a continuous temperature range and extrapolate to generate physically consistent configurations for substantially larger systems (up to $128\times 128$). This capability to learn scale-invariant physical rules from limited data establishes our approach as a computationally superior paradigm for efficiently investigating critical phenomena in the thermodynamic limit.
@@ II Theoretical Structure of Flow Matching
Figure 2: A naive Flow Matching example. When Flow Matching is trained as a deterministic ODE, trajectories from $\mathbf{v_{\theta}}$ deviate from the straight interpolation in Eq. (1). This is because $\mathbf{v_{\theta}}$ learns the expectation of the conditional velocity, $\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)$, induced by the random pairings of $(\mathbf{x}_{0},\mathbf{x}_{1})$ during training (left panel). Consequently, the first sampling step (middle panel) points toward the target mean; with a finite number of iterations, inference preferentially produces samples near this mean, thereby underrepresenting the target variance and fluctuations.
@@ II.1 Probability Flow ODE
Flow Matching admits arbitrary paths, granting flexibility to optimize sampling speed or modeling capacity [25]. In this work, we adopt a simple linear interpolation between base and data for simplicity. For a target data sample $\mathbf{x}_{1}\sim\mathcal{P}_{data}$ and a standard Gaussian noise sample $\mathbf{x}_{0}\sim\mathcal{N}(0,\mathbb{I})$ of the same dimensionality, we define the path $\{\mathbf{x}_{t}\}_{t=0}^{1}$ , where $t$ is a virtual time parameter indexing the continuous evolution of the distribution. The path at each subscripted time $t$ is given by
@@
Here, $p_{t}(\mathbf{x}_{t}|\phi)$ is the conditional probability path. For example, suppose that $\mathcal{P}_{data}$ is the Boltzmann distribution $exp(-\beta H)/Z$. The configuration $\phi$ of conditions is only the spin configurations $\mathbf{x}_{1}$ (generally, the condition can include temperatures, coupling constants, and other information that may affect the target distribution). Consider the path $\mathbf{x}_{t}$ in Eq. (1), because we choose $\mathbf{x}_{0}$ that satisfies the standard Gaussian distribution, therefore, conditional probability $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$ is the Gaussian distribution with mean $t\mathbf{x}_{1}$ and variance $(1-t)^{2}$ [28]. The $p_{t}(\mathbf{x}_{t})$ hence reads as
@@
The corresponding ODE to $\mathbf{u_{t}}(\mathbf{x}_{t})$ is also called the probability flow ODE [26]:
@@
And we want the neural network $\mathbf{v_{\theta}}$ with learnable parameters $\theta$ to learn this velocity field $\mathbf{u_{t}}(\mathbf{x}_{t})$.
@@ II.2 Loss Function
However, we do not know the exact velocity field $\mathbf{u_{t}}(\mathbf{x}_{t})$ from the training data. Thus, use the trick that the gradient of the loss function $\mathcal{L}(\theta)$ to update the neural network $\mathbf{v_{\theta}}$ is only depend on $\theta$, it is equivalent [25] to define conditional Flow Matching loss $\mathcal{L}_{CFM}(\theta)$ that we want to minimize as
@@ II.3 Flow of Random Sampling
Although the path in Eq. (1) is straight, but the training paths are created by randomly pairing data $\mathbf{x}_{1}$ and noise $\mathbf{x}_{0}$ with different conditions. Therefore, different conditional trajectories passing through the same $x_{t}$ at time $t$ generally have different directions, and the velocity field $\mathbf{u_{t}}(\mathbf{x}_{t})$ is not identical to the constant vector $(\mathbf{x}_{1}-\mathbf{x}_{0})$ (see Fig. 2). What the neural network $\mathbf{v_{\theta}}(\mathbf{x}_{t},t|\phi)$ learned is an expectation over conditional velocities and there is inherent uncertainty about the direction. One way to understand this is to view the training path as a quantum process, where a sample from the initial distribution $\mathbf{x}_{0}$ has the opportunity to reach any target distribution $\mathbf{x}_{1}$. Flow matching allows the neural network to learn the generators corresponding to the path that contributes most, the classical path. Due to this reason, instead of collapsing sampling to a single jump from noise to data, we must integrate the learned velocity field over a few time steps (however, research on the one-step sampling is rapidly expanding [29, 30]). Concretely, we divide the interval $t\in[0,1]$ into $N$ steps $\{t_{k}\}^{N}_{k=0}$ and numerically solve the problem.
@@
III Simulation Framework
@@ III.1 Parameter-Wide Sampling via a Single Network
The spin stiffness is then defined as [2]:
@@
Figure 4: Susceptibility $\chi(T)$ for a $32\times 32$ XY lattice. The points show results from MCMC (MC) simulations, which serve as both the ground truth and the training data for our Flow Matching (FM) model. The three panels demonstrate the model’s performance when trained on this MCMC data at varying levels of precision: (a) a dense grid with temperature spacing $\Delta T=0.05$, (b) $\Delta T=0.10$, and (c) a sparse grid with $\Delta T=0.50$. The training data consists of $5000$ configurations per temperature point. In all cases, the FM model is evaluated on a much finer grid ($\Delta T=0.01$), showing it successfully reproduces the BKT-related features even when trained on the sparsest data set.
IV Result I: Accurate Interpolation of Parameter Density
Flow Matching can learn the transfer of the data background distribution through limited data, so the most intuitive benefit is that one can train the model through discrete samples in parameter space and then make continuous predictions. Just like when one pulls the tablecloth, the petals on the tablecloth will move accordingly. When the tablecloth pulling method is learned and fixed, no matter how many petals are sprinkled, there will be the same movement pattern. Here we show that Flow Matching trained with MC data at intervals of 0.1 in the temperature range of $0.1\sim 2.0$ can be easily sampled with ten times the temperature density (intervals of 0.01). Figs. 3 show that the spin configuration sampling results of Flow Matching still retain good physical statistical properties after increasing the temperature density.
Flow Matching provides an additional benefit and advantage in parameter interpolation. Certain physical quantities, such as susceptibility, often exhibit complex and abrupt changes near phase transition points as temperature varies. When data points are sparse along the temperature axis, conventional numerical fitting or machine learning methods typically struggle to perform effective interpolation. However, although these physical observables change abruptly and complexly, the underlying probability distributions of the samples may vary simply and uniformly concerning certain parameters, such as temperature. In the example of the two-dimensional XY model, as long as Flow Matching accurately learns the relationship between the probability distribution of samples and temperature, it can effectively generate samples within temperature ranges where training data are sparse. Fig. 4 shows that Flow Matching, trained on spin configuration samples that are extremely sparse in temperature, can still produce samples that exhibit qualitatively consistent statistical properties.
V Result II: Reliable Extrapolation of System Sizes

### v4

# Efficient Identification of Critical Transitions via Flow Matching: A Scalable Generative Approach for Many-Body Systems

@@ Abstract
We propose a machine learning framework based on Flow Matching (FM) to identify critical properties in many-body systems efficiently. Using the 2D XY model as a benchmark, we demonstrate that a single network, trained only on configurations from a small ($32\times 32$) lattice at sparse temperature points, effectively generalizes across both temperature and system size. This dual generalization enables two primary applications for large-scale computational physics: (i) a rapid ”train-small, predict-large” strategy to locate phase transition points for significantly larger systems ($128\times 128$) without retraining, facilitating efficient finite-size scaling analysis; and (ii) the fast generation of high-fidelity, decorrelated initial spin configurations for large-scale Monte Carlo simulations, providing a robust starting point that bypasses the long thermalization times of traditional samplers. These capabilities arise from the combination of the Flow Matching framework, which learns stable probability-flow vector fields, and the inductive biases of the U-Net architecture that capture scale-invariant local correlations. Our approach offers a scalable and efficient tool for exploring the thermodynamic limit, serving as both a rapid explorer for phase boundaries and a high-performance initializer for high-precision studies.
@@ I Introduction
Monte Carlo—especially Markov chain Monte Carlo (MCMC)—is one of the most powerful computational tools across three fronts: in condensed matter, it is the workhorse for probing phase transitions and complex systems [Landau_Binder_2005]; in quantum many-body physics, quantum Monte Carlo is among the few reliable non-perturbative methods [Sandvik_2010, RevModPhys.73.33]; and in lattice QCD, Euclidean-lattice Monte Carlo provides the only systematically improvable first-principles non-perturbative approach at low energies [RevModPhys.82.1349, PhysRevLett.49.613]. Algorithmically, MCMC targets the Boltzmann measure via detailed balance and ergodicity (e.g. Metropolis, HMC [HMC], cluster [PhysRevLett_Wolff]), with uncertainties governed by the integrated autocorrelation time [Landau_Binder_2005].
@@
However, MCMC encounters critical slowing down as correlation lengths grow near continuous phase transitions and toward the continuum limit; in lattice QCD, topological modes can freeze, inflating autocorrelations and complicating error estimates [Schaefer_2011]. Fermionic models further face the sign problem, whose generic resolution is NP-hard [PhysRevLett.94.170201]. Progress relies on advanced updates [PhysRevLett_Wolff], tempering/multilevel schemes [Hukushima_1996], and ML-assisted samplers; notably, flow-based generative sampling have shown decorrelated proposals and efficiency gains in lattice field theories [PhysRevD_Flow_based_MCMC, kanwar2024flow_based_sampling_lattice_field, abbott2024practical_app_flows, PhysRevD_NF_MC_sampling, PhysRevD.106.014514].
@@
Figure 1: Schematic diagram of our work. Using Flow Matching with U-Net [ronneberger2015unet] inductive biases, we learn a deterministic probability-flow sampler that—trained only on $L_{0}\times L_{0}$ lattices—simultaneously interpolates across temperatures and extrapolates (dashed lines) to physically consistent configurations of $L_{1}\times L_{1}$ lattices with $L_{1}\gg L_{0}$, revealing scale-invariant rules and enabling a fast and stable exploration of critical phenomena near the critical temperature, $T_{c}$. Here $\chi$ denotes the susceptibility, and the green shadow represents the critical regime.
@@
Early successes with models like Normalizing Flows (NFs) demonstrated the potential to eliminate critical slowing down for a single set of system parameters [PhysRevD_Flow_based_MCMC, hackett2025_flow_based_sampling_multimodal_extendedmode]. However, the inefficiency of these initial approaches in exploring parameter space is a significant limitation. Studying critical phenomena requires simulations across various parameters (e.g., temperatures), but training a separate, computationally expensive model for each point is impractical [PhysRevD_NF_MC_sampling]. A key advancement to address this has been the development of conditional models, which learn the system’s behavior across a continuous parameter range from sparse training data, enabling efficient interpolation and extrapolation [PhysRevD_NF_MC_sampling, Gerdes_2023LQF_equi_flow]. These models are often implemented with either architecturally rigid NFs, which can suffer from training instabilities like mode collapse [hackett2025_flow_based_sampling_multimodal_extendedmode, PhysRevDmode_collapse_for_flow_based], or with Diffusion Models (DM) [zhang2025efficient_unbiased_sampling_boltzmann, Wang2024_Diffusion_models_as_stochastic_quantization, Fukushima_2025_SQ_DM, zhu2024diffusionmodelslatticegauge, zhu2025physics_conditioned_DM], which, while robust, typically rely on a computationally intensive iterative sampling process.
@@
Even as conditional models begin to address the challenge of exploring parameter space, a more fundamental and computationally demanding limitation persists in both traditional MCMC and these new generative approaches: the lack of generalizability across system sizes. Systematic studies of critical phenomena and the thermodynamic limit hinge on a sequence of simulations on progressively larger lattices. Under the current paradigm, each lattice size requires an entirely new, independent simulation or model training, a process whose computational cost scales prohibitively. Therefore, a truly efficient framework would not only interpolate across parameters like temperature but also capture the local, scale-independent correlation patterns, enabling extrapolation from a single, small-system training process to much larger lattices. Our work is designed to address this critical gap directly.
@@
In this work, we employ Flow Matching (FM)  [lipman2023flowmatchinggenerativemodeling], a next-generation framework that combines the advantages of these preceding approaches while mitigating their drawbacks. FM directly learns the deterministic vector field of a continuous transformation, known as the probability flow ODE [song2021score_matching]. This provides a more stable regression-based training objective than traditional NFs and enables significantly faster sampling than the stochastic processes of DMs [ddpm, song2021score_matching]. This training stability, architectural flexibility, and sampling efficiency combination makes FM an ideal candidate for building a robust and generalizable conditional sampler. When integrated with the strong inductive biases of a U-Net architecture [ronneberger2015unet], we demonstrate that this framework offers an effective solution to a crucial challenge: achieving a robust dual generalization. We show that a single model, trained only on data from a small lattice ($32\times 32$), can simultaneously interpolate across a continuous temperature range and extrapolate to predict configurations that reflect the essential critical behavior of substantially larger systems (up to $128\times 128$). This capability to generalize local interactions from limited data establishes our approach as a computationally efficient strategy for locating phase transitions in the thermodynamic limit.
@@ II Theoretical Structure of Flow Matching
Figure 2: A naive Flow Matching example. When Flow Matching is trained as a deterministic ODE, trajectories from $\mathbf{v_{\theta}}$ deviate from the straight interpolation in Eq. (1). This is because $\mathbf{v_{\theta}}$ learns the expectation of the conditional velocity, $\mathbf{v_{t}}(\mathbf{x}_{t}|\phi)$, induced by the random pairings of $(\mathbf{x}_{0},\mathbf{x}_{1})$ during training (left panel). Consequently, the first ODE iteration step (middle panel) points toward the target mean; with a finite number of iterations, inference preferentially produces samples near this mean, thereby underrepresenting the target variance and fluctuations.
@@ II.1 Probability Flow ODE
Flow Matching admits arbitrary paths, granting flexibility to optimize sampling speed or modeling capacity [lipman2023flowmatchinggenerativemodeling]. In this work, we adopt a simple linear interpolation between base and data for simplicity. For a target data sample $\mathbf{x}_{1}\sim\mathcal{P}_{data}$ and a standard Gaussian noise sample $\mathbf{x}_{0}\sim\mathcal{N}(0,\mathbb{I})$ of the same dimensionality, we define the path $\{\mathbf{x}_{t}\}_{t=0}^{1}$ , where $t$ is a virtual time parameter indexing the continuous evolution of the distribution. The path at each subscripted time $t$ is given by
@@
Here, $p_{t}(\mathbf{x}_{t}|\phi)$ is the conditional probability path. For example, suppose that $\mathcal{P}_{data}$ is the Boltzmann distribution $exp(-\beta H)/Z$. The configuration $\phi$ of conditions is only the spin configurations $\mathbf{x}_{1}$ (generally, the condition can include temperatures, coupling constants, and other information that may affect the target distribution). Consider the path $\mathbf{x}_{t}$ in Eq. (1), because we choose $\mathbf{x}_{0}$ that satisfies the standard Gaussian distribution, therefore, conditional probability $p_{t}(\mathbf{x}_{t}|\mathbf{x}_{1})$ is the Gaussian distribution with mean $t\mathbf{x}_{1}$ and variance $(1-t)^{2}$ [lipman2024flow_matching_guide_code]. The $p_{t}(\mathbf{x}_{t})$ hence reads as
@@
The corresponding ODE to $\mathbf{u_{t}}(\mathbf{x}_{t})$ is also called the probability flow ODE [song2021score_matching]:
@@
And we want the neural network $\mathbf{v_{\theta}}$ with learnable parameters $\theta$ to learn this velocity field $\mathbf{u_{t}}(\mathbf{x}_{t})$. There are several methods for using neural networks to predict velocity fields. One reliable method is to let the neural network $\mathbf{net_{\theta}}$ predict the target sample and calculate the corresponding velocity field at different time steps [JIT]:
$\mathbf{v_{\theta}}(\mathbf{x}_{t},t)\equiv\frac{\mathbf{net_{\theta}}(\mathbf{x}_{t},t)-\mathbf{x}_{t}}{1-t}$ (9)
Which can be derived from Eq. (1) and Eq. (2).
@@ II.2 Loss Function
However, we do not know the exact velocity field $\mathbf{u_{t}}(\mathbf{x}_{t})$ from the training data. Thus, use the trick that the gradient of the loss function $\mathcal{L}(\theta)$ to update the neural network $\mathbf{v_{\theta}}$ is only depend on learnable parameters $\theta$, it is equivalent [lipman2023flowmatchinggenerativemodeling] to define conditional Flow Matching loss $\mathcal{L}_{CFM}(\theta)$ that we want to minimize as
@@ II.3 Flow of Random Sampling
Although the path in Eq. (1) is straight, but the training paths are created by randomly pairing data $\mathbf{x}_{1}$ and noise $\mathbf{x}_{0}$ with different conditions. Therefore, different conditional trajectories passing through the same $x_{t}$ at time $t$ generally have different directions, and the velocity field $\mathbf{u_{t}}(\mathbf{x}_{t})$ is not identical to the constant vector $(\mathbf{x}_{1}-\mathbf{x}_{0})$ (see Fig. 2). What the neural network $\mathbf{v_{\theta}}(\mathbf{x}_{t},t|\phi)$ learned is an expectation over conditional velocities and there is inherent uncertainty about the direction. One way to understand this is to view the training path as a quantum process, where a sample from the initial distribution $\mathbf{x}_{0}$ has the opportunity to reach any target distribution $\mathbf{x}_{1}$. Flow matching allows the neural network to learn the generators corresponding to the path that contributes most, the classical path. Due to this reason, instead of collapsing sampling to a single jump from noise to data, we must integrate the learned velocity field over a few time steps (however, research on the one-step sampling is rapidly expanding [geng2025meanflowsonestepgenerative, frans2025stepdiffusionshortcutmodels]). Concretely, we divide the interval $t\in[0,1]$ into $N$ steps $\{t_{k}\}^{N}_{k=0}$ and numerically solve the problem (see Fig. 2).
@@
III Generative Framework and Observable Estimation
@@ III.1 Parameter-Wide Sampling via a Single Network
The spin stiffness is then defined as [Sandvik_2010]:
@@
Figure 4: Susceptibility $\chi(T)$ for a $32\times 32$ XY lattice. The points show results from MCMC (MC) simulations, which serve as both the ground truth and the training data for our Flow Matching (FM) model. The three panels demonstrate the model’s performance when trained on this MCMC data at varying levels of precision: (a) a dense grid with temperature spacing $\Delta T=0.05$, (b) $\Delta T=0.10$, and (c) a sparse grid with $\Delta T=0.50$. The training data consists of $5000$ configurations per temperature point. In all cases, the FM model is evaluated on a much finer grid ($\Delta T=0.01$), showing it successfully reproduces the BKT-related features even when trained on the sparsest data set (In (c), Monte Carlo training data is available only at the four temperatures outside the phase transition point.). Since our method directly generates spin configurations rather than predicting physical quantities, it is well-suited for generating high-quality initial spin configurations for Monte Carlo simulations that require high accuracy.
IV Result I: Dense Estimation of Observables via Interpolation
Flow Matching can learn the transfer of the data background distribution through limited data, so the most intuitive benefit is that one can train the model through discrete samples in parameter space and then estimate observables continuously. Just like when one pulls the tablecloth, the petals on the tablecloth will move accordingly. When the tablecloth pulling method is learned and fixed, no matter how many petals are sprinkled, there will be the same movement pattern. Here we show that Flow Matching trained with MC data at intervals of 0.1 in the temperature range of $0.1\sim 2.0$ can be easily sampled with ten times the temperature density (intervals of 0.01). Figs. 3 show that the spin configuration sampling results of Flow Matching still retain good physical observable trends after increasing the temperature density.
Flow Matching provides an additional benefit and advantage in parameter interpolation. Certain physical quantities, such as susceptibility, often exhibit complex and abrupt changes near phase transition points as temperature varies. When data points are sparse along the temperature axis, conventional numerical fitting or machine learning methods typically struggle to perform effective interpolation. However, although these physical observables change abruptly and complexly, the underlying probability distributions of the samples may vary simply and uniformly concerning certain parameters, such as temperature. In the example of the two-dimensional XY model, as long as Flow Matching accurately learns the relationship between the probability distribution of samples and temperature, it can effectively reconstruct the trend of observables within temperature ranges where training data are sparse. Fig. 4 shows that Flow Matching, trained on spin configuration samples that are extremely sparse in temperature, can still produce configurations that preserve the essential features of the phase transition.
V Result II: Extrapolation for Finite-Size Scaling Analysis